1. Suppose you wish to invest N$2,000 today so that you have N$6,000 six years from now. What must the annual interest rate be to achieve your goal if interest is compounded monthly? [4]

2. Assume that the current rate of inflation on housing is estimated at 10% per annum. Lydia's house is currently valued at N$ 450 000.
(a) How much could you expect to sell it for in 5 years? [3]
(b) Is it worthwhile to sell your house now or in five years, justify your answer? [2]

3. You want to buy a cellphone at the end of the year which will cost you N$ 4500. Assume on 1 January 2020 you open up a Unit trust which pays 6.5% interest per month. How much should you invest (starting from January) to yield N$4500 in December 2020? [4]

4. According to a survey carried out in 2019, the biomass of crabs in a particular region was estimated to be 5 350 MT. In 2014, the biomass was estimated to be 3 897 MT. Calculate the annual rate of growth of the crab’s population? [4]
1
Expert's answer
2020-09-22T14:01:42-0400

1) "2000*(1+r)^{72}=6000"

r-interest rate is compounded monthly.

"r=0.0154"

2) a) "450,000*1.1^5=724,729.5"

b) If Lydia will reinvest this money with rate more than inflation, Yes, to sell hous now is worthwhile. Another situation it is not.

3) "x*(1+0.065)^{12}=4500"

"x=2,113.57"

4) "3,897*(1+r)^{5}=5,350"

"r=0.0654"